A +6.00 -μC point charge is moving at a constant 8.00×106 m/s in the + y-direction, relative to a reference frame. At the instant when the point charge is at the origin of this reference frame, what is the magnetic-field vector it produces at the following points.

Part A: x = +.5 m, y = 0 m, z = 0 m

Part B: x = 0 m, y = -.5 m, z = 0 m

Part C: x = 0 m, y = 0 m, z = +.5 m

Part D: x = 0 m, y = -.5 m, z = +.5 m
Answer #1
Concepts and reason

The required concepts to solve these questions are Biot-Savart law, direction of the magnetic field and the general rule of the cross product.

Biot-Savart Law: This law is used to determine the magnitude and direction of the magnetic field produced. Magnetic field is produced by a moving charge.

Calculate the magnetic field by using the Biot-Savart law at different points.

init

Fundamentals

The expression for Biot-Savart law is given by,

B=μ04πqv×r^r2\vec B = \frac{{{\mu _0}}}{{4\pi }}\frac{{q\vec v \times \hat r}}{{{r^2}}}

Here,qqis the charge, rris the magnitude of the displacement vector, μ0{\mu _0}is the permeability and r^\hat r is the unit vector andvv is the velocity.

The expression of the unit vector can be written as,

r^=rr\hat r = \frac{{\vec r}}{{\left| {\vec r} \right|}}

Here, r^\hat ris the unit vector.

(a)

The expression for Biot-Savart law is given by,

B=μ04πqv×r^r2\vec B = \frac{{{\mu _0}}}{{4\pi }}\frac{{q\vec v \times \hat r}}{{{r^2}}}

Substitute rr\frac{{\vec r}}{{\left| {\vec r} \right|}} for r^\hat rin the above equation.

B=μ04πqv×rr3\vec B = \frac{{{\mu _0}}}{{4\pi }}\frac{{q\vec v \times \vec r}}{{{r^3}}} …… (1)

At x=+0.5mx = + 0.5{\rm{ m}}, the displacement vector is,

r=(0.5m)i+0j+0k\vec r = \left( {0.5{\rm{ m}}} \right)\vec i + 0\vec j + 0\vec k

The magnitude of displacement vector is,

r=(0.5m)2+02+02=0.5m\begin{array}{c}\\r = \sqrt {{{\left( {0.5{\rm{ m}}} \right)}^2} + {0^2} + {0^2}} \\\\ = 0.5{\rm{ m }}\\\end{array}

Substitute 4π×107NA24\pi \times {10^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}} for μ0{\mu _0} , 6μC6{\rm{ \mu C}}forqq, (8×106ms1)j\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec jfor v\vec v, 0.5m0.5{\rm{ m }} for rr and (0.5m)i\left( {0.5{\rm{ m}}} \right)\vec i forr\vec r in the equation (1).

B=(4π×107NA24π)(6μC)(1C106μC)((8×106ms1)j×(0.5m)i(0.5m)3)=(107NA2)(6×106C)((4×106m2s2)(k)0.125m3)=1.92×105T(k)=(0T)i+(0T)j(1.92×105T)k\begin{array}{c}\\\vec B = \left( {\frac{{4\pi \times {{10}^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}}}}{{4\pi }}} \right)\left( {6{\rm{ \mu C}}} \right)\left( {\frac{{1{\rm{ C}}}}{{{{10}^6}{\rm{ \mu C}}}}} \right)\left( {\frac{{\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec j \times \left( {0.5{\rm{ m}}} \right)\vec i}}{{\left( {0.5{\rm{ m}}} \right){{\rm{ }}^3}}}} \right)\\\\ = \left( {{{10}^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}}} \right)\left( {6 \times {{10}^{ - 6}}{\rm{ C}}} \right)\left( {\frac{{\left( {4 \times {{10}^6}{\rm{ }}{{\rm{m}}^2} \cdot {{\rm{s}}^{ - 2}}} \right)\left( { - \vec k} \right)}}{{0.125{\rm{ }}{{\rm{m}}^3}}}} \right)\\\\ = 1.92 \times {10^{ - 5}}\,{\rm{T}}\left( { - \vec k} \right)\\\\\, = \left( {0{\rm{ T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j - \left( {1.92 \times {{10}^{ - 5}}\,{\rm{T}}} \right)\vec k\\\end{array}

(b)

Use equation (1) to calculate the magnetic field at point y=0.5my = - 0.5{\rm{ m}}.

B=μ04πqv×rr3\vec B = \frac{{{\mu _0}}}{{4\pi }}\frac{{q\vec v \times \vec r}}{{{r^3}}}

The displacement vector is,

r=0i+(0.5m)j+0k\vec r = 0\vec i + \left( { - 0.5{\rm{ m}}} \right)\vec j + 0\vec k

The magnitude of displacement vector is,

r=02+(0.5m)2+02=0.5m\begin{array}{c}\\r = \sqrt {{0^2} + {{\left( { - 0.5{\rm{ m}}} \right)}^2} + {0^2}} \\\\ = 0.5{\rm{ m}}\\\end{array}

Substitute 4π×107NA24\pi \times {10^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}} for μ0{\mu _0} , 6μC6{\rm{ \mu C}}forqq, (8×106ms1)j\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec jfor v\vec v, 0.5m0.5{\rm{ m }} for rr and (0.5m)j\left( { - 0.5{\rm{ m}}} \right)\vec j forr\vec r in the equation (1).

B=(4π×107NA24π)(6μC)(1C106μC)((8×106ms1)j×(0.5m)j(0.5m)3)=(0T)i+(0T)j+0k\begin{array}{c}\\\vec B = \left( {\frac{{4\pi \times {{10}^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}}}}{{4\pi }}} \right)\left( {6{\rm{ \mu C}}} \right)\left( {\frac{{1{\rm{ C}}}}{{{{10}^6}{\rm{ \mu C}}}}} \right)\left( {\frac{{\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec j \times \left( { - 0.5{\rm{ m}}} \right)\vec j}}{{\left( {0.5{\rm{ m}}} \right){{\rm{ }}^3}}}} \right)\\\\ = \left( {0{\rm{ T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j + 0\vec k\\\end{array}

(c)

Use equation (1) to calculate the magnetic field at point z=0.5mz = 0.5{\rm{ m}}.

B=μ04πqv×rr3\vec B = \frac{{{\mu _0}}}{{4\pi }}\frac{{q\vec v \times \vec r}}{{{r^3}}}

The displacement vector is,

r=0i+0j+(0.5m)k\vec r = 0\vec i + 0\vec j + \left( {0.5{\rm{ m}}} \right)\vec k

The magnitude of displacement vector is,

r=02+02+(0.5m)2=0.5m\begin{array}{c}\\r = \sqrt {{0^2} + {0^2} + {{\left( {0.5{\rm{ m}}} \right)}^2}} \\\\ = 0.5{\rm{ m}}\\\end{array}

Substitute 4π×107NA24\pi \times {10^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}} for μ0{\mu _0} , 6μC6{\rm{ \mu C}}forqq, (8×106ms1)j\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec jfor v\vec v, 0.5m0.5{\rm{ m }} for rr and (0.5m)k\left( {0.5{\rm{ m}}} \right)\vec k forr\vec r in the equation (1).

B=(4π×107NA24π)(6μC)(1C106μC)((8×106ms1)j×(0.5m)k(0.5m)3)=(1.92×105T)i+(0T)j+(0T)k\begin{array}{c}\\\vec B = \left( {\frac{{4\pi \times {{10}^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}}}}{{4\pi }}} \right)\left( {6{\rm{ \mu C}}} \right)\left( {\frac{{1{\rm{ C}}}}{{{{10}^6}{\rm{ \mu C}}}}} \right)\left( {\frac{{\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec j \times \left( {0.5{\rm{ m}}} \right)\vec k}}{{\left( {0.5{\rm{ m}}} \right){{\rm{ }}^3}}}} \right)\\\\ = \left( {1.92 \times {{10}^{ - 5}}\,{\rm{T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j + \left( {0{\rm{ T}}} \right)\vec k\\\end{array}

(d)

Use equation (1) to calculate the magnetic field at points x=0mx = 0{\rm{ m}},y=0.5my = - 0.5{\rm{ m}},z=+0.5mz = + 0.5{\rm{ m}}

B=μ04πqv×rr3\vec B = \frac{{{\mu _0}}}{{4\pi }}\frac{{q\vec v \times \vec r}}{{{r^3}}}

The displacement vector is,

r=0i+(0.5m)j+(0.5m)k\vec r = 0\vec i + \left( { - 0.5{\rm{ m}}} \right)\vec j + \left( {0.5{\rm{ m}}} \right)\vec k

The magnitude of displacement vector is,

r=02+(0.5m)2+(0.5m)2=0.707m\begin{array}{c}\\r = \sqrt {{0^2} + {{\left( {0.5{\rm{ m}}} \right)}^2} + {{\left( {0.5{\rm{ m}}} \right)}^2}} \\\\ = 0.707{\rm{ m}}\\\end{array}

Substitute 4π×107NA24\pi \times {10^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}} for μ0{\mu _0} , 6μC6{\rm{ \mu C}}forqq, (8×106ms1)j\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec jfor v\vec v, 0.707m0.707{\rm{ m}} for rr and 0i+(0.5m)j+(0.5m)k0\vec i + \left( { - 0.5{\rm{ m}}} \right)\vec j + \left( {0.5{\rm{ m}}} \right)\vec k forr\vec r in the equation (1).

B=(4π×107NA24π)(6μC)(1C106μC)((8×106ms1)j×((0.5m)j+(0.5m)k)(0.707m)3)=(107NA2)(6×106C)((4×106m2s1)i0.353m3)=(6.79×106T)i+(0T)j+(0T)k\begin{array}{c}\\\vec B = \left( {\frac{{4\pi \times {{10}^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}}}}{{4\pi }}} \right)\left( {6{\rm{ \mu C}}} \right)\left( {\frac{{1{\rm{ C}}}}{{{{10}^6}{\rm{ \mu C}}}}} \right)\left( {\frac{{\left( {8 \times {{10}^6}{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}} \right)\vec j \times \left( {\left( { - 0.5{\rm{ m}}} \right)\vec j + \left( {0.5{\rm{ m}}} \right)\vec k} \right)}}{{{{\left( {0.707{\rm{ m}}} \right)}^3}}}} \right)\\\\ = \left( {{{10}^{ - 7}}{\rm{ N}} \cdot {{\rm{A}}^{ - 2}}} \right)\left( {6 \times {{10}^{ - 6}}{\rm{ C}}} \right)\left( {\frac{{\left( {4 \times {{10}^6}{{\rm{m}}^2} \cdot {{\rm{s}}^{ - 1}}} \right)\vec i}}{{0.353{\rm{ }}{{\rm{m}}^3}}}} \right)\\\\ = \left( {6.79 \times {{10}^{ - 6}}\,{\rm{T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j + \left( {0{\rm{ T}}} \right)\vec k\\\end{array}

Ans: Part a

The magnetic field at points x=+0.5mx = + 0.5{\rm{ m}}, y=0my = 0{\rm{ m}}, z=0mz = 0{\rm{ m}}are (0T)i+(0T)j(1.92×105T)k\left( {0{\rm{ T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j - \left( {1.92 \times {{10}^{ - 5}}\,{\rm{T}}} \right)\vec k.

Part b

The magnetic field at points x=0mx = 0{\rm{ m}},y=0.5my = - 0.5{\rm{ m}},z=0mz = 0{\rm{ m}}are(0T)i+(0T)j+0k\left( {0{\rm{ T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j + 0\vec k.

Part c

The magnetic field at points x=0mx = 0{\rm{ m}},y=0y = 0,z=0.5mz = 0.5{\rm{ m}}are (1.92×105T)i+(0T)j+(0T)k\left( {1.92 \times {{10}^{ - 5}}\,{\rm{T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j + \left( {0{\rm{ T}}} \right)\vec k.

Part d

The magnetic field at points x=0mx = 0{\rm{ m}},y=0.5my = - 0.5{\rm{ m}},z=+0.5mz = + 0.5{\rm{ m}}are (6.79×106T)i+(0T)j+(0T)k\left( {6.79 \times {{10}^{ - 6}}\,{\rm{T}}} \right)\vec i + \left( {0{\rm{ T}}} \right)\vec j + \left( {0{\rm{ T}}} \right)\vec k.